Name the Property.
a: 6+1*2=1*2+6
b: (7x)*y=7(xy)
c: 0+a=a
d: 2(x+3)=2x+2*3
e: 1/2(8*5)=(1/2*8)*5
f: 1*(-2/3)=-2/3

Answer by stanbon(75887)

Name the Property.
a: 6+1*2=1*2+6
Ans: commutative
-----------------------
b: (7x)*y=7(xy)
Ans: associative
-----------------------
c: 0+a=a
Ans: additive identity
------------------------
d: 2(x+3)=2x+2*3
Ans: distributive
------------------------
e: 1/2(8*5)=(1/2*8)*5
Ans: associative
-----------------------
f: 1*(-2/3)=-2/3
Ans: multiplicative identity
==================================
Cheers,
Stan H.